OpenStudy (anonymous):

If z= 1/(1-cosθ-isinθ) then Re(z)=

OpenStudy (anonymous):

u want to solve it ?

OpenStudy (anonymous):

yes

OpenStudy (koikkara):

\(\color{red}{~COMPLEX~NUMBERS}\) \(\color{darkgreen}{~''Nice~To~Meet~You~!!"}\) @dsfaaa  \(Hmm...\) (1 - cosθ - isinθ) = 2cos^2(θ/2) - 2isin(θ / 2)cos(θ / 2) =2cos(θ / 2)[cos(θ / 2)-isin(θ / 2)] (1-cosθ+isinθ= 2sin^2(θ / 2)+2isin(θ / 2)cos(θ / 2)=2sin(θ / 2)[sin(θ / 2)+icos(θ / 2)] so z=cot(θ /2) [cos(θ / 2)-isin(θ / 2)]/[sin(θ / 2)+icos(θ / 2)] \(Multiply~top~and~bottom~by~'i~' ~\) \(iC\)(c-is)/(-c+is) \(where....\) C=cot(θ / 2), c=cos(θ / 2) and s=sin(θ / 2) This gives \(-iC\), Since c-is=-(-c+is) \(Hope~it~helps~!!\) Don't Forget, \(Appreciate~the~Best~Answer\) with a click on \(\color{purple}{"Best Response"}\) Button Too ! \(\color{darkblue}{~Have~A~Nice~Day~!!}\)

OpenStudy (koikkara):

\(Note:~It's~just~A~Sample,~Showing~how~to~solve~and~its~not~exact~Answer~!!\)

OpenStudy (anonymous):

\[\large \begin{align*}z&=\frac{1}{1-\cos\theta-i\sin\theta}\\\\ &=\frac{1}{1-\cos\theta-i\sin\theta}\cdot\frac{1-\cos\theta+i\sin\theta}{1-\cos\theta+i\sin\theta}\\\\ &=\frac{1-\cos\theta+i\sin\theta}{(1-\cos\theta)^2+\sin^2\theta}\\\\ &=\frac{1-\cos\theta+i\sin\theta}{1-2\cos\theta+\cos^2\theta+\sin^2\theta}\\\\ &=\frac{1-\cos\theta+i\sin\theta}{2-2\cos\theta}\\\\ &=\frac{1-\cos\theta}{2(1-\cos\theta)}+i\frac{\sin\theta}{2-2\cos\theta}\\\\ &=\frac{1}{2}+i\frac{\sin\theta}{2-2\cos\theta} \end{align*}\]